# uniINF: Best-of-Both-Worlds Algorithm for Parameter-Free Heavy-Tailed MABs

### v1

@@ 6 Conclusion
A.1 $z_{t,i}$ is Multiplicatively Close to $x_{t,i}$: Proof of Lemma 4
@@
A.2 Calculating $w_{t,i}$ via Partial Derivatives
@@
A.3 Bregman Divergence before Expectation: Proof of Lemma 5
@@
A.4 Bounding the Expectation of Learning Rate $S_{T+1}$: Proof of Lemma 6
@@
A.5 Adversarial Bounds for Bregman Divergence Terms: Proof of Theorem 7
@@
A.6 Stochastic Bounds for Bregman Divergence Terms: Proof of Theorem 8
@@
Appendix B $\Psi$-Shifting Terms: Omitted Proofs in Section 5.3
B.1 $\Psi$-Shifting Terms before Expectation: Proof of Lemma 9
@@
B.2 Adversarial Bounds for $\Psi$-Shifting Terms: Proof of Theorem 10
In adversarial case, we have already bounded the expectation of $\mathbb{E}[S_{T+1}]$ in Section A.5. Therefore, we have the following lemma.
@@ B.2 Adversarial Bounds for $\Psi$-Shifting Terms: Proof of Theorem 10
B.3 Stochastic Bounds for $\Psi$-Shifting Terms: Proof of Theorem 11
@@
Appendix C Skipping Loss Terms: Omitted Proofs in Section 5.4
C.1 Constant Clipping Threshold Argument: Proof of Lemma 12
@@
C.2 Adversarial Bounds for Skipping Losses: Proof of Theorem 13
@@
C.3 Stochastic Bounds for Skipping Losses: Proof of Theorem 14
@@
Appendix D Main Theorem: Proof of Theorem 3
@@
The following lemma characterizes the FTRL regret decomposition, which is the extension of the classical FTRL bound (Lattimore and Szepesvári,, 2020, Theorem 28.5). Dann et al., 2023b (, Lemma 17) also gave a similar result, but we include a full proof here for the sake of completeness.

### v2

@@ Appendix B Discussions on Assumption 1 (Truncated Non-Negative Loss)
C.1 $z_{t,i}$ is Multiplicatively Close to $x_{t,i}$: Proof of Lemma 4
@@
C.2 Calculating $w_{t,i}$ via Partial Derivatives
@@
C.3 Bregman Divergence before Expectation: Proof of Lemma 5
@@
C.4 Bounding the Expectation of Learning Rate $S_{T+1}$: Proof of Lemma 6
@@
C.5 Adversarial Bounds for Bregman Divergence Terms: Proof of Theorem 7
@@
C.6 Stochastic Bounds for Bregman Divergence Terms: Proof of Theorem 8
@@
Appendix D $\Psi$-Shifting Terms: Omitted Proofs in Section 5.3
D.1 $\Psi$-Shifting Terms before Expectation: Proof of Lemma 9
@@
D.2 Adversarial Bounds for $\Psi$-Shifting Terms: Proof of Theorem 10
In adversarial case, we have already bounded the expectation of $\mathbb{E}[S_{T+1}]$ in Section C.5. Therefore, we have the following lemma.
@@ D.2 Adversarial Bounds for $\Psi$-Shifting Terms: Proof of Theorem 10
D.3 Stochastic Bounds for $\Psi$-Shifting Terms: Proof of Theorem 11
@@
Appendix E Skipping Loss Terms: Omitted Proofs in Section 5.4
E.1 Constant Clipping Threshold Argument: Proof of Lemma 12
@@
E.2 Adversarial Bounds for Skipping Losses: Proof of Theorem 13
@@
E.3 Stochastic Bounds for Skipping Losses: Proof of Theorem 14
@@
Appendix F Main Theorem: Proof of Theorem 3
@@
The following lemma characterizes the FTRL regret decomposition, which is the extension of the classical FTRL bound (Lattimore & Szepesvári, 2020, Theorem 28.5). Dann et al. (2023a, Lemma 17) also gave a similar result, but we include a full proof here for the sake of completeness.